OpenStudy (anonymous):

Find the zeros of the function. f(x) = 3x3 - 12x2 - 15x

OpenStudy (tkhunny):

Factor factor factor. Let's see what you get.

OpenStudy (anonymous):

take 3x common \[\Large 3x(x^2-4x-5)=0\] \[\Large 3x=0\] \[\Large x^2-4x-5=0\] use quadratic formula for \[\Large x^2-4x-5=0\]

OpenStudy (anonymous):

@nezzy can you do this now?